OpenStudy (anonymous):

A ball is rolling down a hill. The ball accelerates from rest to 2 m.s in 4 seconds. a)What is the acceleration? b)How far does the ball roll in the 4 seconds? please show the work too. im so confused on this.

OpenStudy (agreene):

average acceleration is computed by: \[\frac{\Delta \alpha}{\Delta t}\] or the change in accel divided by change in t. In this case: (2-0)/(4-0) 2/4=1/2 <= average acceleration with part b, think of it as this: 4s*1/2m/s^2 = 2 m/s *4 s = 8m if I'm remembering that bit right O.o I really want to use calculus because it makes more sense